# How to Leverage Diverse Demonstrations in Offline Imitation Learning

### v2

@@ 5 Experiments
Baselines. We evaluate our method against six strong baseline methods in offline IL: 1) BCE, the standard BC trained only on expert demonstrations; 2) BCU, BC trained on union data; 3) DWBC (Xu et al., 2022), an offline IL method that leverages suboptimal demonstrations by jointly training a discriminator to re-weight the BC objective; 4) ISWBC (Li et al., 2023), an offline IL method that adopts importance sampling to enhance BC; 5) CSIL (Watson et al., 2023), a model-free IRL method that learns a shaped reward function using the BC policy; 6) MLIRL (Zeng et al., 2023), a model-based offline IRL algorithm based on bi-level optimization.
@@ A.1 Benchmarks
We evaluate our method on a number of environments (Robomimic, MuJoCo, Adroit, FrankaKitchen, and AntMaze) which are widely used in prior studies (Nakamoto et al., 2023; Watson et al., 2023). We elaborate in what follows.
@@ A.3 Baselines
We evaluate our method against six strong baseline methods in offline IL: 1) Behavior Cloning with Expert Data (BCE), the standard BC trained only on expert demonstrations; 2) Behavior Cloning with Union Data (BCU), BC trained on union data; 3) Discriminator-Weighted Behavioral Cloning (DWBC) (Xu et al., 2022), an offline IL method that leverages suboptimal demonstrations by jointly training a discriminator to re-weight the BC objective (https://github.com/ryanxhr/DWBC); 4) Importance-Sampling-Weighted Behavioral Cloning (ISWBC) (Li et al., 2023), an offline IL method that adopts importance sampling to enhance BC (https://github.com/liziniu/ISWBC); 5) Coherent Soft Imitation Learning (CSIL) (Watson et al., 2023), a model-free IRL method that learns a shaped reward function by entropy-regularized BC (https://joemwatson.github.io/csil); 6) Maximum Likelihood-Inverse Reinforcement Learning (MLIRL) (Zeng et al., 2023), a recent model-based offline IRL algorithm based on bi-level optimization (https://github.com/Cloud0723/Offline-MLIRL).

### v3

@@ 5 Experiments
Baselines. We evaluate our method against six strong baseline methods in offline IL: 1) BCE, the standard BC trained only on expert demonstrations; 2) BCU, BC trained on union data; 3) DWBC (Xu et al., 2022), an offline IL method that leverages suboptimal demonstrations by jointly training a discriminator to re-weight the BC objective; 4) ISWBC (Li et al., 2023), an offline IL method that adopts importance sampling to enhance BC; 5) CSIL (Watson et al., 2024), a model-free IRL method that learns a shaped reward function using the BC policy; 6) MLIRL (Zeng et al., 2023), a model-based offline IRL algorithm based on bi-level optimization.
@@ A.1 Benchmarks
We evaluate our method on a number of environments (Robomimic, MuJoCo, Adroit, FrankaKitchen, and AntMaze) which are widely used in prior studies (Nakamoto et al., 2023; Watson et al., 2024). We elaborate in what follows.
@@ A.3 Baselines
We evaluate our method against six strong baseline methods in offline IL: 1) Behavior Cloning with Expert Data (BCE), the standard BC trained only on expert demonstrations; 2) Behavior Cloning with Union Data (BCU), BC trained on union data; 3) Discriminator-Weighted Behavioral Cloning (DWBC) (Xu et al., 2022), an offline IL method that leverages suboptimal demonstrations by jointly training a discriminator to re-weight the BC objective (https://github.com/ryanxhr/DWBC); 4) Importance-Sampling-Weighted Behavioral Cloning (ISWBC) (Li et al., 2023), an offline IL method that adopts importance sampling to enhance BC (https://github.com/liziniu/ISWBC); 5) Coherent Soft Imitation Learning (CSIL) (Watson et al., 2024), a model-free IRL method that learns a shaped reward function by entropy-regularized BC (https://joemwatson.github.io/csil); 6) Maximum Likelihood-Inverse Reinforcement Learning (MLIRL) (Zeng et al., 2023), a recent model-based offline IRL algorithm based on bi-level optimization (https://github.com/Cloud0723/Offline-MLIRL).